OpenStudy (anonymous):

the measure of angle A is 38 degrees less than its compliement

OpenStudy (anonymous):

let angle A = x deg then its compliment angle = x + 38 x + (x + 38) = 90 2x = 90-38 x = 52/2 x = 26 so angleA is 26deg its complement is 64 deg